OpenStudy (solomonzelman):

Just a couple seconds tutoring on limits, something I got confused about.

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \lim_{x \rightarrow ∞} ~~\frac{2x^2-5}{x(x+4)} } \] I saw the answer was 2, but didn't quite get why. I foiled the bottom and got \[\huge\color{blue}{ \lim_{x \rightarrow ∞} ~~\frac{2x^2-5}{x^2+4x} } \] then using L'H'S I get \[\huge\color{blue}{ \lim_{x \rightarrow ∞} ~~\frac{(2x^2-5)'}{(x^2+4x)'} } \] \[\huge\color{blue}{ \lim_{x \rightarrow ∞} ~~\frac{4x-0}{2x+4} } \] \[\huge\color{blue}{ \lim_{x \rightarrow ∞} ~~\frac{2x}{x+2} } \] so the final answer would then be 0/2 or zero, not 2, right?

OpenStudy (anonymous):

no it would be 2 , \[\frac{ (2x)' }{ (x+2)' }=\frac{ 2 }{ 1 }\]

OpenStudy (solomonzelman):

deriving it a second time, after having done it once already?

OpenStudy (anonymous):

yeah , as long as you have limit of form inf/inf you can do it

OpenStudy (solomonzelman):

So doing it the standard way, deriving once and plugging in 0, is not correct? I thought L'H'S is derive top and bottom and then plug in whatever x is approaching to.

OpenStudy (solomonzelman):

I never thought then in inf limit you can derive it more than once.

OpenStudy (anonymous):

why you plug in 0 if limit goes to infinity

OpenStudy (solomonzelman):

oh, yeah, didn't think abt that, my bad; ty

OpenStudy (solomonzelman):

I see, so using L'H'S in inf limits I'll just keep deriving until I get a constant (or a number without a variable)

OpenStudy (anonymous):

yeah u can derive as many time as you want until the condition of L'H'S is satisfied

OpenStudy (solomonzelman):

Ok, ty!

OpenStudy (anonymous):

ur welcome

OpenStudy (isaiah.feynman):

Review L'Hospital's rule once more.

OpenStudy (solomonzelman):

I am pretty sure I did that just now, ty for your suggestion, Isaiah.